# IMU: Influence-guided Machine Unlearning

### v2

@@ 4.1 Setup
Note: IU is the improved version of IU presented in Jia et al. (2023), and SCAR denotes the version using ${\mathcal{D}}_{f}$ only.
@@
Table 1: Quantitative results on CIFAR-10 and CIFAR-100. Results are averaged over all 10 classes for class-wise unlearning on CIFAR-10, across 20 superclasses, and across 5 subclasses in one superclass for subclass-wise unlearning on CIFAR-100.
Dataset and models. We evaluate MU methods on image classification using CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009) datasets with the model ResNet-18 (He et al. 2016). Person re-identification, which matches a person’s identity across different cameras or locations in a video or image sequence, is also included. ResNet-50 with a fully-connected layer is trained on the Market-1501 dataset (Zheng et al. 2015), which contains labeled person IDs captured under varying camera views. We also assess unlearning in a sequence modeling setting using a GPT-2 model on synthetic data (Fan et al. 2024a). For large language model unlearning, we employ the Llama-3.2-3B-Instruct model on the TOFU benchmark (Dorna et al. 2025).
Scenarios. We benchmark across five different unlearning settings: (i) Class-wise unlearning: forget a full category from CIFAR-10. (ii) Subclass-wise unlearning: forget a subclass within a superclass (e.g., boy from people) in CIFAR-100. (iii) Sample-wise unlearning (see appendix): randomly select samples from ${\mathcal{D}}_{\texttt{train}}$ as the forget set. (iv) Person re-id unlearning: forget a specific identity from the 751 IDs in the Market-1501 train set, and evaluate model generalization on the query and gallery sets. (v) Distributional unlearning in sequence modeling: forget two Markov sub-distributions in a synthetic task. (vi) LLM unlearning (see appendix): perform the forget05 task on the TOFU benchmark.
Baselines. (1) Retrain, (2) Gradient Ascent (GA) (Thudi et al. 2022), (3) Random Label (RL) (Jia et al. 2023), (4) Influence Unlearning (IU) (Jia et al. 2023), (5) SCAR (Bonato, Cotogni, and Sabetta 2024), without the information about ${\mathcal{D}}_{r}$, (6) SSD (Foster, Schoepf, and Brintrup 2024), (7) NPO (Zhang et al. 2024), (8) SimNPO (Fan et al. 2024a).
Metrics. (1) Accuracy on ${\mathcal{D}}_{f}$, ${\mathcal{D}}_{r}$, and ${\mathcal{D}}_{t}$, denoted as $\texttt{Acc}_{{\mathcal{D}}_{f}}$, $\texttt{Acc}_{{\mathcal{D}}_{r}}$, and $\texttt{Acc}_{{\mathcal{D}}_{t}}$, respectively. (2) Membership inference attacks (MIA) aim to infer information about the training data. (3) $W_{\texttt{dist}}$ (Tarun et al. 2023), i.e., Wasserstein-1 distance, by measuring the similarity between the output distributions of the unlearned model and the retrained model on ${\mathcal{D}}_{r}$. (4) mAP (mean average precision) represents the average of the area under the precision-recall curve for each query, reflecting the overall retrieval quality.
4.2 Results on class-wise unlearning
We first evaluate on CIFAR-10, trying to forget an entire class category. As shown in Table 1, our proposed method IMU achieves superior unlearning performance. IU fails to effectively preserve the model utility, showing significant drops of $\sim$15% in accuracy in both ${\mathcal{D}}_{r}$ and the test set ${\mathcal{D}}_{t}$. GA and RL perform even worse in this regard. The SOTA retain-data-free MU method SCAR demonstrates strong forgetting ability; however, it comes at the cost of notable degradation in model utility, with accuracies on ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{t}$ reduced by approximately 14%. We further employ the potent LLM unlearning method, NPO, for image classification. It significantly outperforms baselines in both the forget quality and model utility preservation.
@@ 4.2 Results on class-wise unlearning
4.3 Results on subclass-wise unlearning
We further evaluate MU methods on CIFAR-100, trying to forget a single subclass from a semantic superclass. Specifically, (1) for subclass-wise unlearning across all super-classes, we randomly select a subclass from each of the 20 super-classes in CIFAR-100 to be forgotten, and report the average performance; (2) for subclass-wise unlearning within a single super-class, we focus on the super-class ‘people’, where each time, one of the five subclasses is selected to be forgotten while the remaining four are retained. These settings are more challenging than the standard class-wise unlearning, as subclasses within a superclass share semantically and visually similar features, making the forgetting more fine-grained and less separable.
Despite the increased difficulty, IMU consistently demonstrates better performance than other MU methods. As shown in Table 1, all MU methods completely remove the knowledge about ${\mathcal{D}}_{f}$, while IMU best preserving the performance over ${\mathcal{D}}_{r}$, having an accuracy around $98\%$ and achieves a high test accuracy compared to other baselines.
Method mAP ($\uparrow$) Top-1 ($\uparrow$) Top-5 ($\uparrow$) Run time ($\downarrow$) Original 68.50 85.45 94.27 IU 44.79 69.69 84.59 147 GA 04.12 18.59 32.66 043 RL 00.83 00.02 00.06 104 SSD 51.79 73.57 87.35 137 NPO 41.40 63.13 81.03 055 SCAR 50.97 73.87 88.54 135 IMU (Ours) 55.85 76.07 88.75 043
Table 2: Unlearning on person re-identification.
Method Model utility Forget quality
$\ell_{r}$ ($\downarrow$) $\ell_{r}^{\texttt{KL}}$ ($\downarrow$) $\ell_{f}$ ($\uparrow$) $\ell_{f}^{\texttt{KL}}$ ($\uparrow$)
Original 1.99 - 2.18 -
GA 4.24 2.21 6.95 3.42
NPO 3.97 1.95 7.26 3.46
SimNPO 4.61 2.57 7.44 3.83
IMU (Ours) 3.86 1.83 7.53 3.96
Table 3: Unlearning on a sequence modeling problem.
4.4 Results on person re-identification
Pedestrians’ body shapes, clothing, facial features, and other biometric characteristics are often sensitive and uniquely identifiable, making them important targets for protection in MU scenarios. To this end, we also evaluate MU methods on the person re-identification task. This task involves matching images of the same individual captured under varying camera viewpoints or multiple cameras. Here, we aim to forget all samples associated with a particular identity (e.g., $\texttt{pid}=1$). This task poses unique challenges to MU methods, as forgetting a specific individual requires precise removal of identity-related features while preserving general person-level recognition performance.
We adopt GradCAM (Selvaraju et al. 2017) to visualize regions where models focus on w/o and w/ IMU. As shown in Figure 3, when evaluated on ${\mathcal{D}}_{r}$, IMU preserves focus on discriminative yet identity-agnostic regions. In contrast, when evaluated on ${\mathcal{D}}_{f}$, our scrubbed model significantly shifts attention away from identity-revealing regions (e.g., clothing logos, face, and hairstyle), indicating successful removal of sensitive cues. In Table 2, we further evaluate model utility by measuring mAP and Top-$k$ accuracy on a query-gallery split consisting of individuals entirely unseen during training. In general, IMU achieves a good trade-off between forget quality and model utility.
Figure 3: Visualizations of regions where models focus on for ${\mathcal{D}}_{f}$ and ${\mathcal{D}}_{r}$, respectively. For each triplet, from left to right are the original image, the activation map generated by the original model, and IMU scrubbed model, respectively.
4.5 Case study: Markov chains
Aside from vision tasks, we further evaluate our method on a simple sequence modeling problem. Following the setting in SimNPO (Fan et al. 2024a), we construct a mixture of Markov chains with a state space of size 10, the retain distribution consists of Markov chains that transition uniformly among states $\{1,2,3\}$, while the forget distribution is a mixture of two Markov chains Forget1 (transition uniformly among $\{4,5,6\}$) and Forget2 (transition uniformly among $\{7,8,9\}$) with equal probability. Each sequence has length $T$ and is denoted as $s=(s_{1},s_{2},\cdots,s_{T})$ where $s_{t}\in\{0,1,\cdots,9\}$ for $t\in[0,T]$. A GPT-2 model is trained to learn the conditional distribution. Details can be found in §7 of SimNPO (Fan et al. 2024a). Table 3 presents results compared to NPO and SimNPO. Performance is evaluated by the loss values over the retain and forget sets, as well as the respective KL divergence distance calculated compared with the retrained model. Our method surpasses NPO and SimNPO in both forget quality and model utility.
4.6 Ablation study
@@
Table 4: Impact of varying the ratio ($r$) of top-ranked ${\mathcal{D}}_{f}$ (selected based on ${\mathcal{I}}^{c}({\bm{x}},{\mathcal{D}}_{f})$) on unlearning performance.
@@
Table 5: Impact of frequency ($\nu$) of influence value updates on unlearning performance. $\nu=0$ means only update at the first epoch; $\nu=1$ and $\nu=2$ mean update every epoch and every two epochs, respectively.
@@
We finally conduct ablation studies to analyze the effects of two key design choices in our method: (1) the ratio $r$ of top-ranked forgetting data selected based on influence values, and (2) the update frequency $v$ of influence scores during the unlearning process. Table 4 presents the effect of varying the top-$r$ fraction of the forgetting data ${\mathcal{D}}_{f}$ for unlearning. As $r$ decreases, the model retains strong forget quality as evidenced by near-zero accuracy on ${\mathcal{D}}_{f}$, while achieving slightly improved accuracy on ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{t}$. Notably, using only 5% of the most influential data points is sufficient to induce forgetting, while also enhancing generalization and utility, highlighting the efficiency of targeting high-impact data points. Table 5 investigates the impact of varying the frequency $v$ of influence score updates. All variants achieve strong forget quality, with minor differences in model utility. Infrequent updates are sufficient for maintaining effective unlearning, suggesting that IMU is robust.
5 Conclusion, Limitations, Broader Impacts

### v3

@@ 3.3. Validity of Classifier-level Approximation
Remark 3.4.
In practice, $\Delta_{-{\bm{x}}_{i}}\ell({\mathcal{D}}_{f})$ can be heavy-tailed due to estimation noise and sample heterogeneity. Using $\psi(w)=w$ may lead to a few extreme samples dominating the weighted objective in Equation 5. We therefore introduce a smoothing transform $\psi(\cdot)$ to control the weight concentration. We consider several monotone choices in our experiments, including $\psi(w)=\sqrt{w}$ and $\psi(w)=\tanh(w)$, as well as a softmax-based normalization $\psi(w)=\exp(w)$. Comparison can be found in Table 5 in the appendix.
Remark 3.5.
Early MU methods often rely on full-parameter updates, which provide a generic solution but become impractical for large-scale models due to high compute/memory cost and unnecessary interference with retained knowledge. Recent work, therefore, moves toward selective, parameter-efficient unlearning that updates only a small subset of parameters (e.g., SalUn (Fan et al., 2024b) and LoRA-style approaches such as LUNE (Liu et al., 2025)) which inject task-specific incremental parameters without modifying the base model. Following this paradigm, IMU performs head-only unlearning, enabling efficient forgetting with minimal utility degradation.
Figure 3. Divergence comparison between GA, NPO, and IMU.
3.4. Connection with NPO
Negative Preference Optimization (NPO) (Zhang et al., 2024) stands out in LLM unlearning, it defines
(11) ${\mathcal{L}}_{\texttt{NPO}}({\bm{\theta}})\coloneqq\frac{2}{\beta}\mathbb{E}_{{\mathcal{D}}_{f}}\left[\log\left(1+\left(\frac{\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})}{\pi_{\texttt{ref}}({\bm{y}}|{\bm{x}})}\right)^{\beta}\right)\right],$
where $\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})$ denotes the output probability of the current model, and $\pi_{\texttt{ref}}({\bm{y}}|{\bm{x}})$ is the output probability from a reference model. For image classification, the model typically outputs a softmax probability distribution over classes as $\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})=\texttt{softmax}\left(f_{{\bm{\theta}}}({\bm{x}})\right)$, where $f_{{\bm{\theta}}}({\bm{x}})$ are the logits, and $\ell({\bm{x}},{\bm{y}};{\bm{\theta}})=-\log\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})$.
The gradients of GA, NPO, and IMU evaluated at ${\bm{x}}_{i}$ are
(12) ${\bm{g}}_{\texttt{GA}}=-\nabla_{{\bm{\theta}}}\ell({\bm{x}}_{i},{\bm{y}}_{i};{\bm{\theta}}),$
(13) ${\bm{g}}_{\texttt{NPO}}=-w_{\texttt{NPO},i}\cdot\nabla_{{\bm{\theta}}}\ell({\bm{x}}_{i},{\bm{y}}_{i};{\bm{\theta}}),$
(14) ${\bm{g}}_{\texttt{IMU}}=-\hat{w}_{i}\cdot\nabla_{{\bm{\theta}}}\ell({\bm{x}}_{i},{\bm{y}}_{i};{\bm{\theta}}),$
where $w_{\texttt{NPO},i}=2\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})^{\beta}/\left(\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})^{\beta}+\pi_{\texttt{ref}}({\bm{y}}|{\bm{x}})^{\beta}\right)$ is the adaptive smoothing weight, controlling the divergence speed. NPO adaptively assigns weights to data points to help prevent catastrophic collapse in GA; yet, Fan et al. (2024a) shows that NPO exhibits reference model bias, causing an uneven allocation of unlearning power. IMU instead uses explicitly estimated influence scores to guide unlearning. In this sense, IMU can be viewed as an influence-weighted counterpart to GA, offering a more direct and interpretable mechanism for prioritizing which data points to forget.
Theorem 3.6 (Divergence Analysis).
Consider logistic regression with labels $y\in\{0,1\}$ and the observed-label likelihood $\pi_{{\bm{\theta}}}(y\mid{\bm{x}})\triangleq\sigma\big((2y-1)z\big)$ where $z={\bm{\theta}}^{\top}{\bm{x}}+b$ and $\sigma(u)=\frac{1}{1+\exp(-u)}$. Assume the model is trained on $\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}$ and denote ${\bm{X}}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{n}]^{\top}\in\mathbb{R}^{n\times d}$, ${\bm{K}}={\bm{X}}{\bm{X}}^{\top}\in\mathbb{R}^{n\times n}$ be the Gram matrix. Then, after $t$ SGD updates with step size $\eta$, the following divergence bounds hold:
$\|{\bm{\theta}}^{t}_{\texttt{GA}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\eta^{2}\lambda_{\max}({\bm{K}})^{2}nt^{2},$
$\|{\bm{\theta}}^{t}_{\texttt{NPO}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\eta^{2}\lambda_{\max}({\bm{K}})^{2}nt^{2}\|{\bm{W}}_{\texttt{NPO}}\|^{2},$
(15) $\|{\bm{\theta}}^{t}_{\texttt{IMU}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\eta^{2}\lambda_{\max}({\bm{K}})^{2}nt^{2}\|{\bm{W}}_{\texttt{IMU}}\|^{2}.$
where $\lambda_{\max}({\bm{K}})$ denote the largest eigenvalues of $\bm{K}$}, and ${\bm{W}}_{\texttt{NPO}}\triangleq\mathrm{diag}(w_{\texttt{NPO},1},\ldots,w_{\texttt{NPO},n}),{\bm{W}}_{\texttt{IMU}}\triangleq\mathrm{diag}(\hat{w}_{1},\ldots,\hat{w}_{n})$.
The proof of Theorem 3.6 can be found in § A.2. Theorem 3.6 provides a coarse worst-case upper bound on the parameter drift under the weighted norm $\|\cdot\|_{{\bm{X}}^{\top}{\bm{X}}}$. In particular, GA corresponds to uniform weighting, while NPO and IMU introduce per-sample weights, which can be viewed as a form of data importance weighting that reallocates unlearning strength across samples. From this perspective, NPO can be interpreted as a confidence-based attribution (relative to the reference model), whereas IMU uses influence estimates to attribute which forgetting data are most supportive of fitting ${\mathcal{D}}_{f}$ and prioritizes their removal. Note that this result is not intended to establish a strict ordering of divergence speeds among GA, NPO, and IMU, but rather to make explicit how per-sample weighting enters the drift behavior. We therefore complement it with empirical comparisons under tuned learning rates.
Specifically, Figure 3 presents the class-wise forgetting on CIFAR-10. We observe that NPO reaches low forget accuracy (around 2.37%) quickly and then stabilizes, but it is accompanied by a continuous decline in model utility, from 92.01% to 85.89%. In contrast, IMU exhibits a more gradual decrease in forget accuracy while consistently maintaining a higher test accuracy around 92.60%. Further comparisons on language tasks are reported in § 4 and § B.
@@ 3.4. Connection with NPO
Table 1. Quantitative results on CIFAR-10 and CIFAR-100. Results are averaged over all 10 classes for class-wise unlearning on CIFAR-10, across 20 superclasses, and across 5 subclasses in one superclass for subclass-wise unlearning on CIFAR-100.
@@
Note: IU is an improved version presented in the paper (Fan et al., 2024b), SCAR denotes the version using ${\mathcal{D}}_{f}$ only.
4. Experiment
4.1. Setup
Dataset & models. We evaluate on image classification using CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009) with ResNet-18 (He et al., 2016). Person re-identification, which matches a person’s identity across different cameras or locations in a video or image sequence, is also included. ResNet-50 with a fully-connected layer is trained on the Market-1501 dataset (Zheng et al., 2015), which contains labeled person IDs captured under varying camera views. We also assess MU in a sequence modeling setting using a GPT-2 model on synthetic data (Fan et al., 2024a). For LLM unlearning, we employ the Llama-3.2-3B-Instruct model on TOFU (Dorna et al., 2025).
Scenarios. We benchmark across five different unlearning settings: (i) Class-wise unlearning: forget a full category from CIFAR-10. (ii) Subclass-wise unlearning: forget a subclass within a superclass (e.g., boy from people) in CIFAR-100. (iii) Sample-wise unlearning (see appendix): randomly select samples from ${\mathcal{D}}_{\texttt{train}}$ as the forget set. (iv) Person re-id unlearning: forget a specific identity from the 751 IDs in the Market-1501 train set, and evaluate model generalization on the query and gallery sets. (v) Distributional unlearning in sequence modeling (see appendix): forget two Markov sub-distributions in a synthetic task. (vi) LLM unlearning (see appendix): perform forget05 task on TOFU.
Baselines. (1) Retrain, (2) Gradient Ascent (GA) (Thudi et al., 2022), (3) Random Label (RL) (Golatkar et al., 2020a), (4) Influence Unlearning (IU(Fan et al., 2024b), (5) SCAR (Bonato et al., 2024), (6) SSD (Foster et al., 2024a), (7) NPO (Zhang et al., 2024), (8) SimNPO (Fan et al., 2024a).
Metrics. (1) Accuracy on ${\mathcal{D}}_{f}$, ${\mathcal{D}}_{r}$, and ${\mathcal{D}}_{t}$, denoted as $\texttt{Acc}_{{\mathcal{D}}_{f}}$, $\texttt{Acc}_{{\mathcal{D}}_{r}}$, and $\texttt{Acc}_{{\mathcal{D}}_{t}}$, respectively. (2) Membership inference attacks (MIA(Fan et al., 2024b) aim to infer information about the training data. (3) $W_{\texttt{dist}}$ (Tarun et al., 2023), i.e., Wasserstein-1 distance, by measuring the similarity between the output distributions of the unlearned model and the retrained model on ${\mathcal{D}}_{r}$. (4) mAP (mean average precision) represents the average of the area under the precision-recall curve for each query, reflecting the overall retrieval quality.
4.2. Main results
Results on class-wise unlearning. We first evaluate on CIFAR-10, trying to forget an entire class category. As shown in Table 1, our proposed method IMU achieves superior unlearning performance.
IU fails to effectively preserve the model utility, showing significant drops of $\sim$15% in accuracy in both ${\mathcal{D}}_{r}$ and the test set ${\mathcal{D}}_{t}$. GA and RL perform even worse in this regard. The SOTA retain-data-free MU method SCAR demonstrates strong forgetting ability; however, it comes at the cost of notable degradation in model utility, with accuracies on ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{t}$ reduced by approximately 14%. We further employ the potent LLM unlearning method, NPO, for image classification. It significantly outperforms baselines in both the forget quality and model utility preservation.
@@ 4.2. Main results
Results on subclass-wise unlearning. We further evaluate MU methods on CIFAR-100, trying to forget a single subclass from a semantic superclass. Specifically, (1) for subclass-wise unlearning across all super-classes, we randomly select a subclass from each of the 20 super-classes in CIFAR-100 to be forgotten, and report the average performance; (2) for subclass-wise unlearning within a single super-class, we focus on the super-class ‘people’, where each time, one of the five subclasses is selected to be forgotten while the remaining four are retained. These settings are more challenging than the standard class-wise unlearning, as subclasses within a superclass share semantically and visually similar features, making the forgetting more fine-grained and less separable.
Despite the increased difficulty, IMU consistently demonstrates better performance than other methods. In Table 1, all MU methods completely remove the knowledge about ${\mathcal{D}}_{f}$, while IMU best preserving the performance over ${\mathcal{D}}_{r}$, having an accuracy around $98\%$ and achieves a high test accuracy compared to other baselines.
Table 2. Unlearning on person re-identification.
@@
Results on person re-identification. Pedestrians’ body shapes, clothing, facial features, and other biometric characteristics are often sensitive and uniquely identifiable, making them important targets for protection in MU scenarios. To this end, we also evaluate MU methods on the person re-identification task. This task involves matching images of the same individual captured under varying camera viewpoints or multiple cameras. Here, we aim to forget all samples associated with a particular identity (e.g., $\texttt{pid}=1$). This task poses unique challenges to MU methods, as forgetting a specific individual requires precise removal of identity-related features while preserving general person-level recognition performance. We adopt GradCAM (Selvaraju et al., 2017) to visualize regions where models focus on w/o and w/ IMU. As shown in Figure 4, when evaluated on ${\mathcal{D}}_{r}$, IMU preserves focus on discriminative yet identity-agnostic regions. In contrast, when evaluated on ${\mathcal{D}}_{f}$, our scrubbed model significantly shifts attention away from identity-revealing regions (e.g., clothing logos, face, and hairstyle), indicating successful removal of sensitive cues. In Table 2, we further evaluate model utility by measuring mAP and Top-$k$ accuracy on a query-gallery split consisting of individuals entirely unseen during training. In general, IMU achieves a good trade-off between forget quality and model utility.
Figure 4. Visualizations of regions where models focus on for ${\mathcal{D}}_{f}$ and ${\mathcal{D}}_{r}$, respectively. For each triplet, from left to right are the original image, the activation map generated by the original model, and IMU scrubbed model, respectively.
Computational complexity analysis. Also note that, for a linear softmax classifier trained with cross-entropy, the Hessian w.r.t. the head parameters admits a closed form as
(16) ${\bm{H}}_{{\bm{\theta}}_{c}^{\ast}}=\frac{1}{|{\mathcal{D}}_{f}|}\sum_{i=1}^{|{\mathcal{D}}_{f}|}{\bm{J}}_{i}^{\top}\left[\operatorname{diag}({\bm{p}}_{i})-{\bm{p}}_{i}{\bm{p}}_{i}^{\top}\right]{\bm{J}}_{i},$
where ${\bm{p}}_{i}=\mathrm{softmax}(f_{{\bm{\theta}}_{c}^{\ast}}({\bm{x}}_{i}))$ denotes the predicted probability vector and ${\bm{J}}_{i}=\frac{\partial f_{{\bm{\theta}}_{c}^{\ast}}({\bm{x}}_{i})}{\partial{\bm{\theta}}_{c}}$ is the Jacobian of the last classifier head. In practice, to mitigate runtime and memory demands, we adopt a Fisher diagonal approximation in implementation and compute it only for the head (size $d$) over $N$ forgetting data every epoch, so the extra overhead of IMU over GA is $\mathcal{O}(NdC)$ where $C$ is the number of classes. Note that all run times reported in our tables include the full pipeline end-to-end, encompassing influence score computation, sample selection and reweighting, as well as the gradient ascent unlearning updates. In experiments, IMU requires on average around twice the run time of GA under the same number of training epochs, indicating that the additional computational cost introduced by the influence estimation stage remains within an acceptable range.
4.3. Ablation study
We finally conduct ablation studies in the class-wise unlearning scenario on CIFAR-10 to analyze the effects of key design choices in our method, including (1) the ratio $r$ of top-ranked forgetting data points selected based on the influence values, (2) the update frequency $\nu$ of influence scores during the unlearning process, (3) the smoothing strategy applied to the influence ${\mathcal{I}}^{c}({\bm{x}}_{i},{\mathcal{D}}_{f})$, (4) the dataset used for influence estimation, and (5) the unfreeze strategy of model layers, as well as the choice of layer for computing influence values. Ablations (3)-(5) can be found in § B in the appendix.
@@ 4.3. Ablation study
Table 3 presents the effect of varying the top-$r$ fraction of the forgetting data ${\mathcal{D}}_{f}$ for unlearning. As $r$ decreases, the model retains strong forget quality as evidenced by near-zero accuracy on ${\mathcal{D}}_{f}$, while achieving slightly improved accuracy on ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{t}$. Notably, using only 5% of the most influential data points is sufficient to induce forgetting, while also enhancing generalization and utility, highlighting the efficiency of targeting high-impact data points.
Table 4 investigates the impact of varying the frequency $\nu$ of influence score updates. All variants achieve strong forget quality, with minor differences in model utility. Infrequent updates are sufficient for maintaining effective unlearning, indicating that the approximation error introduced by using fixed influence weights remains negligible even as the feature space evolves. To maintain fast execution time, we set frequency $\nu=0$ in main experiments.
Additional results provided in the appendix show that different smoothing strategies for ${\mathcal{I}}^{c}({\bm{x}}_{i},{\mathcal{D}}_{f})$ (e.g., sqrt, softmax, tanh) yield similar outcomes, with sqrt performing slightly better overall. Estimating the influence value using only ${\mathcal{D}}_{f}$ is sufficient compared to using ${\mathcal{D}}_{train}$. Moreover, updating only the last classifier layer achieves the best balance between forgetting and preservation, while unfreezing earlier layers or computing influence across all layers significantly degrades performance and increases runtime.
Table 3. Effect of varying the ratio ($r$) of top-ranked ${\mathcal{D}}_{f}$ (selected based on ${\mathcal{I}}^{c}({\bm{x}},{\mathcal{D}}_{f})$) on unlearning performance.
[…]
Table 4. Effect of frequency ($\nu$) of influence value updates on unlearning performance. $\nu=0$ means only update at the first epoch; $\nu=1$ and $\nu=2$ mean update every epoch and every two epochs, respectively.
[…]
5. Conclusion and Limitations